# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v2

@@ Abstract
diversity , ensemble learning , multiple hypotheses prediction , portfolio optimization , predict-then-optimize , structured models
@@ 1 Introduction
The problem of portfolio allocation with a quantitative approach was introduced by Harry Markowitz in 1952 in the so-called Modern Portfolio Theory (MPT) (Markowitz, 1952). It consists of two phases: asset selection and portfolio weight optimization. The asset selection phase decides which assets from a larger set, such as a market, will be included in the portfolio, which has a fixed number of assets but not predetermined choices or weights. The portfolio weight optimization phase determines the percentage of capital allocated to each selected asset. It also has the merit of introducing the concept of diversification as a fundamental component in risk management for investments. In this specific case, diversification is induced by the covariance matrix (Markowitz, 1952). Since then, a large number of methods known as ”plug-in” (or data-driven) approaches have been developed which use optimization inputs with improved estimates of future returns (Best and Grauer, 1991; Broadie, 1993; Giglio et al., 2022).
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.1 Preliminary
Let $\boldsymbol{\hat{w}}$ be an estimator of $\boldsymbol{w}\in\mathcal{R}^{M}$ based on an i.i.d. sample from a distribution $\mathcal{D}$. The prediction error can be decomposed into bias and variance components. In this case, the bias is minimized subject to the portfolio return being above the target value, and the variance is minimized. The optimization problem can be formulated as follows.
$\min_{\boldsymbol{w}}\quad\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}+\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ (1)
$\quad\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$
@@
where $\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-R_{T}\right)^{2}$ represents the bias of the expected return of the portfolio compared to the target return, and $\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ represents the variance of the portfolio. $\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq R_{T}$ ensures that the expected return of the portfolio is above the target return. This setup allows to find the optimal weights $\boldsymbol{\hat{w}}$ that minimize the total prediction error of the portfolio, considering both bias and variance, while ensuring that the expected return of the portfolio is above the target value. The expression in (1) generalizes the standard MSE (when $\boldsymbol{A}=\boldsymbol{I}$), reflecting the bias-variance trade-off (Cai et al., 2024):
$MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})=\mathcal{L}_{\boldsymbol{A}}(\mathbb{E}[\boldsymbol{\hat{w}}]-\boldsymbol{w})+\text{tr}(\boldsymbol{A}\text{Var}[\boldsymbol{\hat{w}}])$ (2)
$\boldsymbol{A}\in\mathcal{R}^{M\times M}$ is a constant positive-definite scaling matrix, $\mathcal{L}_{\boldsymbol{A}}(\boldsymbol{x})=\boldsymbol{x}^{T}\boldsymbol{A}\boldsymbol{x}$ is the $\boldsymbol{A}-norm$ of the vector $\boldsymbol{x}$, and the expectation is evaluated under the distribution $\mathcal{D}$. The first term in (2) vanishes for an unbiased estimator, and the second term vanishes if the estimator has zero variance. Now, consider $\boldsymbol{w}$ as the solution to a decision problem:
$\boldsymbol{w}=\arg\max_{\boldsymbol{w}\in C}P(\boldsymbol{w})$ (3)
for some performance measure $P$ and constraint set $C$. The decision loss for $\boldsymbol{\hat{w}}\in C$ is
$D_{P}(\boldsymbol{\hat{w}})=P(\boldsymbol{w})-\mathbb{E}[P(\boldsymbol{\hat{w}})]\geq 0,$ (4)
where the expectation is taken under $D$. For a given $\boldsymbol{\hat{w}}$ and matrix $\boldsymbol{A}$, establish a relationship between $MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})$ and $D_{P}(\boldsymbol{\hat{w}})$, expressed as a function $f$:
$MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}})=f(D_{P}(\boldsymbol{\hat{w}}))$ (5)
For a parameterized estimator $\boldsymbol{\hat{w}}(\delta)$, compare $\delta_{MSE}=\arg\min_{\delta}MSE_{\boldsymbol{A}}(\boldsymbol{\hat{w}}(\delta))$ with $\delta_{D}=\arg\min_{\delta}D_{P}(\boldsymbol{\hat{w}}(\delta))$ and assess the decision loss increase:
$\Delta_{P}(\boldsymbol{\hat{w}}(\delta))=D_{P}(\boldsymbol{\hat{w}}(\delta_{MSE}))-D_{P}(\boldsymbol{\hat{w}}(\delta_{D}))=\mathbb{E}[P(\boldsymbol{\hat{w}}(\delta_{D}))]-\mathbb{E}[P(\boldsymbol{\hat{w}}(\delta_{MSE}))]\geq 0$ (6)
If $f$ increases and $\Delta_{P}$ is small, the MSE serves as an effective loss function for the decision problem, although these conditions do not imply each other.
In this work, structured ensemble multi-hypothesis prediction models are applied to portfolio optimization. Next, the Portfolio Structured Ensemble Model (PSEM) is introduced, with a loss function based on the generalized bias-variance-diversity decomposition for ensembles (Wood et al., 2024).
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
In this portfolio structured ensemble model (PSEM) framework, the goal is to make portfolio allocation decisions at time $t$ for an investment horizon $\Delta t=T-t$. The daily time-series returns of a portfolio for the out-of-sample holding period $\Delta t$ are denoted as $\boldsymbol{R}_{T}=\{R_{Ti}\}_{i=t}^{T}$. To approximate the expected return, a structured ensemble model predicts the portfolio time-series returns as $\boldsymbol{\hat{R}}_{T}$. This model consists of $M$ individual predictors, $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, where each predictor focuses on forecasting the return of the $j$-th component of the portfolio $\{\boldsymbol{R}_{j}\}_{j=1}^{M}$. The predictors are trained using stochastic gradient descent (SGD). During each training iteration $i$, for a given input $\{\boldsymbol{x}_{ij}\}_{i=1,j=1}^{N,M}$ and its corresponding target $R_{ij}$, the parameters of the $j$-th predictor are updated as:
@@
$\boldsymbol{\theta}_{j}=\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}),R_{ij})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta\left(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))\right),$
where $\eta_{j}$ is the learning rate, and the loss function $\mathcal{L}$ combines a squared error term with L2-regularization and parameter $\lambda_{p}$,
@@
$\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}),R_{ij})=\|f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})-R_{ij}\|_{2}^{2}+\frac{\lambda_{p}}{2N}\sum_{j=1}^{M}\|\boldsymbol{\theta}_{j}\|_{2}^{2}$. The indicator function $\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})))$, which adjusts the extent to which each predictor’s parameters are updated, is defined as:
@@
$\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij})))=\begin{cases}1-\varepsilon,&\text{if }R_{ij}\in\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$
@@
Here, $\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{ij}))$ indicates whether the $j$-th predictor’s output is the top prediction for the $i$-th instance. A smaller diversity parameter $\varepsilon\in[0,1]$, places more emphasis on the top-performing predictors, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble. Once training is complete, the predictors can be used to generate a structured data set from the initial set of inputs, or the predictions obtained during training can be used directly as the structured data set. The predictions of all predictors are organized into a structured data set:
@@
$\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)=\begin{bmatrix}f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{11})&\cdots&f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{1M})\\ \vdots&\ddots&\vdots\\ f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{N1})&\cdots&f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{NM})\end{bmatrix}$ (7)
This structured data set serves as input to a structured ensemble model for prediction. The ensemble output is computed as:
$\boldsymbol{\hat{R}}_{t}=\mathcal{F}(\boldsymbol{\hat{R}}(\varepsilon))\boldsymbol{w}$ (8)
@@
where $\mathcal{F}$ is a set of specific functions of the ensemble model, $\boldsymbol{w}=\{w_{j}\}_{j=1}^{M}$ represents the optimal parameters of the ensemble, which, when properly scaled, are equivalent to portfolio weights. $\boldsymbol{\hat{R}}_{t}$ represents the predicted time series returns for the portfolio in the training data. If $\boldsymbol{\hat{R}}(\varepsilon)$ is computed from the data in period $\Delta t$, the solution is $\boldsymbol{\hat{R}}_{T}$. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
@@
Figure 1 shows a sketch of the model architecture, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $(\varepsilon)$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the s-RBFN shown in this figure and trained using least squares, described in detail in Section 3.3 (Dominguez et al., 2025).
@@
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3 (Dominguez et al., 2025). $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{i})$ is the $j^{th}$ shooter (predictor) with input data $\boldsymbol{x}_{i}$ (for example wind, distance, etc.) and $\boldsymbol{\theta}_{j}$ her shooting skills, $\hat{R}_{ij}$ is the $i^{th}$ shoot of $j^{th}$ shooter and $\boldsymbol{\hat{R}}=\boldsymbol{\hat{R}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{R}_{t}$ is the portfolio target for the training set (equal-weighted portfolio), and $\boldsymbol{w}$ is the s-RBFN model parameters optimized by least-squares.
3.2.1 Diversity in Ensemble Learning
In this section, the Bias-Variance-Diversity decomposition (Wood et al., 2024), is formulated and applied to the problem of portfolio optimization using multiple hypothesis structured prediction models.
If a training set $\{(\boldsymbol{R}_{i},R_{ti})\}_{i=1}^{N}$ drawn from a random variable $D\sim P(\boldsymbol{R},\boldsymbol{R}_{t})^{N}$, with $\boldsymbol{R}_{i}=\{R_{\bullet j}\}_{j=1}^{M}$, and $\boldsymbol{R}=\left[\begin{matrix}R_{11}&\ldots&R_{1M}\\ \vdots&\ddots&\vdots\\ R_{N1}&\ldots&R_{NM}\\ \end{matrix}\right]$. From this, a portfolio ensemble model is learned with predictions, $\boldsymbol{\hat{R}_{T}}$. The expected squared risk can be decomposed as follows:
$\mathbb{E}_{D}\left[\mathbb{E}_{\hat{R}_{t},R_{t}}\left[(\boldsymbol{R}_{t}-\boldsymbol{\hat{R}}_{t})^{2}\right]\right]=$
$\underbrace{\mathbb{E}_{R}\left[\mathbb{E}_{R_{t}|\hat{R}_{t}}\left[(\boldsymbol{R}_{t}-\boldsymbol{R}_{t}^{*})^{2}\right]\right]}_{\text{Noise}}+\underbrace{\mathbb{E}_{\hat{R}_{t}}\left[(\mathbb{E}_{D}[\boldsymbol{\hat{R}}_{t}]-\boldsymbol{R}_{t}^{*})^{2}\right]}_{\text{Bias}}+$
$\underbrace{\mathbb{E}_{\hat{R}_{t}}\left[\mathbb{E}_{D}\left[(\boldsymbol{\hat{R}}_{t}-\mathbb{E}_{D}[\boldsymbol{\hat{R}}_{t}])^{2}\right]\right]}_{\text{Variance}}$ (9)
In this context, $\boldsymbol{R}_{t}^{*}=\mathbb{E}_{R_{t}|R}[\boldsymbol{R}_{t}]$ represent the Bayes-optimal predictions. The first term accounts for the irreducible noise in the problem, which is independent of any model parameters. The second term is the bias, defined as the loss of the expected model relative to the Bayes-optimal predictions. Finally, the third term is the variance, which captures the variation in $\boldsymbol{\hat{R}}_{t}$ due to different training sets.
If $\boldsymbol{R}_{j}$ are the predictions of the $j$-th model in the portfolio ensemble, $\overset{\circ}{\boldsymbol{R}}_{j}$ is the centroid prediction of the $j$-th model in the portfolio ensemble. Assuming a bias-variance decomposition holds in the form of (9), the Generalized Bias-Variance-Diversity decomposition also holds (Wood et al., 2024):
$\mathbb{E}_{D}\left[\mathbb{E}_{\hat{R}_{t},R_{t}}[\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{\hat{R}}_{t})]\right]=\mathbb{E}_{\hat{R}_{t}}\left[\mathbb{E}_{R_{t}|\hat{R}_{t}}[\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{t}^{*})]\right]+\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t}^{*},\overset{\circ}{\boldsymbol{R}}_{j})+$ (10)
$\quad+\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{D}\left[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})\right]-\mathbb{E}_{D}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}}_{t},\boldsymbol{R}_{j})\right]$
where $\overset{\circ}{\boldsymbol{R}}_{j}=\arg\min\mathbb{E}_{D}[\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j})]$ and the formula for the estimated target portfolio is in this case the arithmetic combiner,
$\boldsymbol{\hat{R}}_{t}=\arg\min\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{R}_{j})$, the equal-weighted-portfolio ($1/N$-portfolio rule (DeMiguel et al., 2009)). This decomposition separates the expected ensemble risk into: the noise, which represents the irreducible error inherent in the problem; the average bias of the portfolio returns predictions $\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{R}_{t},\overset{\circ}{\boldsymbol{R}}_{j})$; the average variance of the individual models’ predictions in the portfolio ensemble $\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}[\mathcal{L}(\overset{\circ}{\boldsymbol{R}}_{j},\boldsymbol{R}_{j})]$; the expected ambiguity term, considered as the ensemble diversity, which has the opposite sign to the other terms, $\mathbb{E}\left[\frac{1}{M}\sum_{j=1}^{M}\mathcal{L}(\boldsymbol{\hat{R}}_{t},\boldsymbol{R}_{j})\right]$, capturing the disagreement among the portfolio ensemble members in their predictions.
The Bias-Variance-Diversity decomposition using the MSE, $\mathcal{L}(\boldsymbol{R}_{t},\boldsymbol{\hat{R}}_{t})=(\boldsymbol{R}_{t}-\boldsymbol{\hat{R}}_{t})^{2}$, which implies $\boldsymbol{\hat{R}}_{t}=\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{R}_{j}$, and $\overset{\circ}{\boldsymbol{{R}}_{t}}=\mathbb{E}_{D}[\boldsymbol{\hat{R}}_{t}]$, at each data set $(\boldsymbol{R},\boldsymbol{R}_{t})$ is:
$\mathbb{E}_{D}\left[(\boldsymbol{\hat{R}}_{t}-\boldsymbol{R}_{t})^{2}\right]=$
$=\underbrace{\frac{1}{M}\sum_{j=1}^{M}(\overset{\circ}{\boldsymbol{{R}}_{j}}-\boldsymbol{R}_{t})^{2}}_{\text{Average Bias}}+\underbrace{\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}_{D}\left[(\boldsymbol{{R}}_{j}-\overset{\circ}{\boldsymbol{{R}}_{j}})^{2}\right]}_{\text{Average Variance}}-\underbrace{\mathbb{E}_{D}\left[\frac{1}{M}\sum_{j=1}^{M}(\boldsymbol{{R}}_{j}-\boldsymbol{\hat{R}}_{t})^{2}\right]}_{\text{Diversity}}$ (11)

### v3

@@ Abstract
diversity , ensemble learning , multiple hypotheses prediction , predict-then-optimize , robust portfolio optimization
@@ 1 Introduction
The problem of portfolio allocation with a quantitative approach was introduced by Harry Markowitz in 1952 in the so-called Modern Portfolio Theory (MPT) (Markowitz, 1952). It consists of two phases: asset selection and portfolio weight optimization. The asset selection phase decides which assets from a larger set, such as a market, will be included in the portfolio, which has a fixed number of assets but not predetermined choices or weights. The portfolio weight optimization phase determines the percentage of capital allocated to each selected asset. It also has the merit of introducing the concept of diversification as a fundamental component in risk management for investments. In this specific case, diversification is induced by the covariance matrix (Markowitz, 1952). Since then, a large number of methods known as ”plug-in” (or data-driven) approaches have been developed which use optimization inputs with improved estimates of future returns (Best and Grauer, 1991; Broadie, 1993; Giglio et al., 2022; Kelly and Xiu, 2023).
@@ 2.1 Data-Driven and Robust Optimization
Optimization under uncertainty is critical in many applications, and two prominent approaches are data-driven optimization and robust optimization. Data-driven optimization relies on data to estimate uncertainties, using statistical models or machine learning to incorporate these estimates into decision-making (Bertsimas et al., 2013; Shapiro et al., 2009). It aims to optimize expected performance metrics such as cost or profit by leveraging historical or real-time data. On the other hand, robust optimization focuses on worst-case scenarios, ensuring decisions remain feasible and effective for all outcomes within a predefined uncertainty set, such as polyhedral or ellipsoidal bounds (Ben-Tal et al., 2009). This approach prioritizes stability and reliability, often at the expense of conservatism if the uncertainty set is poorly calibrated (Bertsimas et al., 2011). Recent advances combine the strengths of both approaches, such as distributionally robust optimization (DRO), which optimizes for the worst-case distribution within plausible distributions, bridging probabilistic and robust methodologies (Esfahani and Kuhn, 2015; Rahimian and Mehrotra, 2022). In the context of portfolio optimization, data-driven approaches use historical and real-time market data to model uncertainties, applying methods such as stochastic programming, machine learning, or empirical scenario generation to predict asset returns and covariance structures (Roncalli, 2013; Giglio et al., 2022; Kelly and Xiu, 2023). Building on these, this work presents a mixed method that combines both approaches, featuring a plug-in facet but robust in that the risk diversification level can be adjusted parametrically before optimization, thereby addressing the disconnection between out-of-sample diversification and input data in the optimization process.
@@ 3.1 Preliminary
Given $\boldsymbol{x}\in\mathbb{R}^{N\times M}$ as a vector of returns with $M\geq 2$ assets (each with N time-stamps) having mean $\boldsymbol{\mu}$ and a positive-definite covariance matrix $\boldsymbol{\Sigma}$, then the optimal portfolio $\boldsymbol{w}\in\mathbb{R}^{M}$ can be found by solving the following optimization problem (Markowitz, 1952):
$\min_{\boldsymbol{w}}\quad\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-r\right)^{2}+\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ (1)
$\quad\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq r$
@@
where $\left(\boldsymbol{w}^{\top}\boldsymbol{\mu}-r\right)^{2}$ represents the bias of the expected return of the portfolio $\boldsymbol{w}^{\top}\boldsymbol{\mu}$ compared to the target return $r$, and $\boldsymbol{w}^{\top}\Sigma\boldsymbol{w}$ represents the variance of the portfolio. The constraint $\boldsymbol{w}^{\top}\boldsymbol{\mu}\geq r$ ensures that the expected return of the portfolio is above the target return. The expression in (1) generalizes the standard MSE reflecting the bias-variance trade-off (Cai et al., 2024).
If the aforementioned portfolio optimization can be formulated as a prediction task, then the objective utility function of the mean variance in (1) can be represented as an MSE loss function (Cai et al., 2024). This setup allows for finding the optimal weights that minimize the total prediction error of the portfolio, accounting for both bias and variance, while ensuring that the expected return remains above the target value. In order to incorporate the diversity into such an optimization, this work applies ensemble models to harness diversity in ensemble learning and linking it to portfolio diversification. In a multiple hypothesis prediction setup in which each hypothesis is focused on a particular portfolio constituent, the portfolio can be model as an ensemble of hypotheses so that the diversity term in the Bias-Variance-Diversity decomposition with the MSE Loss can be connected to diversification and controlled parametrically, while the centroid combiner rule is respected in the portfolio target selection. To elaborate, in a supervised setting with the data set of time series returns ($\boldsymbol{x}$, $\boldsymbol{r}$), the Bias-variance-diversity decomposition can be expressed in terms of the MSE loss $\mathcal{L}(\boldsymbol{r},\boldsymbol{\bar{r}})$ as (Wood et al., 2024):
$\underbrace{\frac{1}{M}\sum_{j=1}^{M}(\overset{\circ}{f_{j}}(\boldsymbol{x}_{j})-\boldsymbol{r})^{2}}_{\text{Average Bias}}+\underbrace{\frac{1}{M}\sum_{j=1}^{M}\mathbb{E}\left[(f_{j}(\boldsymbol{x}_{j})-\overset{\circ}{f_{j}}(\boldsymbol{x}_{j}))^{2}\right]}_{\text{Average Variance}}-\underbrace{\mathbb{E}\left[\frac{1}{M}\sum_{j=1}^{M}(f_{j}(\boldsymbol{x}_{j})-\boldsymbol{\bar{r}})^{2}\right]}_{\text{Diversity}}$ (2)
where $f_{j}(\boldsymbol{x}_{j})$ denote the temporal predictions and $\overset{\circ}{f_{j}}(\boldsymbol{x}_{j})=\mathbb{E}[f_{j}(\boldsymbol{x}_{j})]$ represents the centroid prediction of the $j$-th model in the portfolio ensemble respectively, while $\boldsymbol{\bar{r}}=\frac{1}{M}\sum_{j=1}^{M}f_{j}(\boldsymbol{x}_{j})$ represents the centroid combiner rule.
In the next section, the portfolio-structured ensemble model (PSEM) is introduced, where each individual hypothesis predicts a specific portfolio constituent, and the diversity in (2) is linked to portfolio generalization performance and risk diversification. The ensemble targets an equal-weighted portfolio, forming a canonical basis of hypotheses or constituent returns, ensuring that the centroid combiner rule for MSE loss is satisfied.
@@ 3.2 Proposed Portfolio Structured Ensemble Model (PSEM)
The PSEM can be defined as a map $\mathcal{E}:\mathcal{X}\to\boldsymbol{r}$, which aggregates predictions as
$\mathcal{E}(\boldsymbol{x})=g\left(f_{\boldsymbol{\theta}_{1}}(\boldsymbol{x}_{1}),f_{\boldsymbol{\theta}_{2}}(\boldsymbol{x}_{2}),\dots,f_{\boldsymbol{\theta}_{M}}(\boldsymbol{x}_{M})\right)$
from multiple individual predictors or hypotheses $f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})$, parameterized by $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{j}\}_{j=1}^{M}$, to produce portfolio predictions $\boldsymbol{\hat{r}}$. Here, $\mathcal{X}$ is the input space, $\boldsymbol{r}$ is the output space, $f_{\boldsymbol{\theta}_{j}}:\mathcal{X}\to\mathcal{Y}$ represents the $j$-th predictor focused on forecasting the time series returns of the $j$-th component of the portfolio, $\boldsymbol{x}_{j}$, and $g:\mathcal{Y}^{M}\to\boldsymbol{r}$ is the aggregation function (e.g., averaging, voting, or weighted combination). Individual predictors are trained with stochastic gradient descent (SGD) with a parameter $\varepsilon\in[0,1]$ that controls their diversity in learning at each step $i$ so that, for the $j$-th predictor:
@@
$\boldsymbol{\theta}_{j}=\boldsymbol{\theta}_{j}-\eta_{j}\left(\frac{\partial\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(x_{ij}),y_{ij})}{\partial\boldsymbol{\theta}_{j}}+\frac{\lambda_{p}}{N}\boldsymbol{\theta}_{j}\right)\delta\left(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})\right)$
@@
where $y_{ij}=x_{(i+\tau)j}$ with time lag $\tau>0$ is the target, $\eta_{j}$ is the learning rate, and the loss function $\mathcal{L}(f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j}),\boldsymbol{y}_{j})$ combines a squared error term and regularization with parameter $\lambda_{p}$. The indicator function $\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})))$, which adjusts the extent to which each predictor’s parameters are updated, is defined as:
@@
$\delta(\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij})))=\begin{cases}1-\varepsilon,&\text{if }y_{ij}\in\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij}))\\[10.0pt] \frac{\varepsilon}{M-1}&\text{otherwise}\end{cases}$
@@
Here, $\mathcal{Y}_{j}(f_{\boldsymbol{\theta}_{j}}(x_{ij}))$ indicates whether the $j$-th predictor’s output is the top prediction for the $i$-th instance. A smaller $\varepsilon$ places more emphasis on the top-performing predictors, while a larger $\varepsilon$ encourages updates to less accurate predictors, fostering greater diversity in the ensemble.
After the training, the predictions of all predictors form a structured data set:
@@
$\boldsymbol{\hat{x}}^{T}=\boldsymbol{\hat{x}}(\varepsilon)^{T}=\begin{bmatrix}f_{\boldsymbol{\theta}_{1}}(x_{11})&\cdots&f_{\boldsymbol{\theta}_{M}}(x_{1M})\\ \vdots&\ddots&\vdots\\ f_{\boldsymbol{\theta}_{1}}(x_{N1})&\cdots&f_{\boldsymbol{\theta}_{M}}(x_{NM})\end{bmatrix}$ (3)
which serves as input to a structured ensemble model. The PSEM predictions are given by:
$\boldsymbol{\hat{r}}=g(\boldsymbol{\hat{x}}(\varepsilon)^{T})\boldsymbol{w}$ (4)
@@
with ensemble parameters $\boldsymbol{w}$, which, when properly scaled, are equivalent to portfolio weights. By incorporating the diversity parameter $\varepsilon$, the ensemble achieves a balance between prediction accuracy and diversity, enhancing its generalization capabilities in portfolio forecasting tasks. Interestingly, this diversity aligns with the concept of portfolio diversification, as originally proposed in Modern Portfolio Theory (Markowitz, 1952).
@@
Figure 1 shows a sketch of the model architecture, illustrated as a shooting gallery where the shooters represent individual predictors. Each shooter receives inputs such as wind conditions, distance, and target position, aiming to hit their respective targets. In each iteration, the trainer penalizes the best shooter by limiting how much their parameters update during gradient descent, while the remaining shooters update equally. This process introduces diversity in learning through $(\varepsilon)$. Once training is complete, all shots are gathered into a diversified data set, which can be used as input for structured prediction models, like the s-RBFN shown in this figure and trained using least squares, described in detail next (Dominguez et al., 2025).
@@
Figure 1: Structured data set formation, including diversity ($\varepsilon$), and analytical ensemble training of the s-RBFN model, described in Section 3.3. $f_{\boldsymbol{\theta}_{j}}(x_{ij})$ is the $j^{th}$ shooter (predictor) with input data $x_{ij}$ (for example wind, distance, etc.) and $\boldsymbol{\theta}_{j}$ her shooting skills, $\hat{x}_{ij}$ is the $i^{th}$ shoot from $j^{th}$ shooter and $\boldsymbol{\hat{x}}=\boldsymbol{\hat{x}}(\varepsilon)$ is the full set of shoots after training, $\boldsymbol{r}$ is the portfolio target for the training set (equal-weighted portfolio), and $\boldsymbol{w}$ are the s-RBFN model parameters optimized by least-squares.
In the next section, the PSEM optimization in the case of an s-RBFN model is described in detail for obtaining the portfolio weights.